# Biomedical Image Segmentation: A Systematic Literature Review of Deep Learning Based Object Detection Methods

### v1

@@ Abstract
Biomedical image segmentation plays a vital role in diagnosis of diseases across various organs. Deep learning-based object detection methods are commonly used for such segmentation. There exists an extensive research in this topic. However, there is no standard review on this topic. Existing surveys often lack a standardized approach or focus on broader segmentation techniques. In this paper, we conducted a systematic literature review (SLR), collected and analysed 148 articles that explore deep learning object detection methods for biomedical image segmentation. We critically analyzed these methods, identified the key challenges, and discussed the future directions. From the selected articles we extracted the results including the deep learning models, targeted imaging modalities, targeted diseases, and the metrics for the analysis of the methods. The results have been presented in tabular and/or charted forms. The results are presented in three major categories including two stage detection models, one stage detection models and point-based detection models. Each article is individually analyzed along with its pros and cons. Finally, we discuss open challenges, potential benefits, and future research directions. This SLR aims to provide the research community with a quick yet deeper understanding of these segmentation models, ultimately facilitating the development of more powerful solutions for biomedical image analysis.
@@ 1 Introduction
Biomedical image segmentation plays a vital role in the analysis and diagnosis of diseases related to various organs. In this regard, the accurate identification of the region of interest in the biomedical images is a significant phase for identifying affected regions in the image. There are several alternative methods to achieve this segmentation. object detection based methods is a type of instance segmentation that is very common in this domain. Deep learning-based object detection methods offer significant advantages over the other alternatives. Considering the prominence of deep learning based object detection-based methods for biomedical image segmentation, we believe a comprehensive review of this field can be a a valuable work to provide a quick insight into the field. Although, there are few reviews (section 1.2) in the related topic, there is no standard review on the target direction.
@@ 1.1 Background and Motivation
In the instance segmentation, each pixel of image is assigned to some specific instance which has a significant role in biomedical imaging for the identification and analysis of various types of diseases or same type of disease with different locations in the image. Different types of algorithms are used for instance segmentation e.g., supervised learning (SL) models and semi-supervised learning (SSL) approaches. Each approach has its own drawbacks and benefits associated with them. Recently, fully supervised models have got proper attention for biomedical image instance segmentation which are based on human experts’ knowledge for annotations. These models are very laborious and expensive in terms of obtaining dense and accurate annotations  [9]. To overcome the problem of time consumption and human efforts associated with annotation for supervised learning approaches, some semi-supervised models have been proposed which require small set of annotated biomedical image samples as compared to the fully supervised instance segmentation models  [10]. The instance segmentation is mainly categorized into two main types namely proposal-free  [11] and proposal-based instance segmentation  [12]. The former category is based on morphological properties and instance-aware features whereas the latter is type is based on object detection approach which is further divided into two stage detectors (2SD), one stage detectors (1SD), and point-based detectors (PBD) or advanced detectors.
@@
Table 1: COMPARISON WITH RELATED SURVEYS
@@
S.No Authors Covered up to Year Imaging Modalities Processing Tasks Performed Models Targeted Diseases/Organs Covered
@@
1 Isin, Ali et al.  [8] 2016 MRI (T1, T1-Gd, T2, FLAIR) Automatic, Semantic Segmentation CNN, ML, Generative Models Brain Tumor
@@
2 Anwar, Syed Muhammad et al.  [13] 2018 CT, MRI, PET, X-Rays, Ultrasound Segmentation, Classification, Diagnosis CNN different variants + ML models Lungs, Breast, Brain, Thyroid
@@
3 Haque and Neubert  [14] 2019 Clinical, X-Rays, CT, MRI, Ultrasound, OCT, Microscopic Images Semantic Segmentation, Classification, Analysis ML, CNN, RBM, Auto encoders, GANs, RNN, FCN, U-Net, ResNet, V-Net Brain, Breast, Liver, Spleen, Heart, Stomach, Ventricles, Arteries, Skin
@@
4 Yahya and Boubakeur  [15] 2019 PET, CT, X-Ray, Ultrasound, Histological Images Classification, Segmentation VGG-16, GoogleNet, AlexNet, ResNet, ResUnet, SegNet, UNet++, RD-Unet, RNN, ML Models Not Mentioned
@@
5 Tongxue Zhou et al.  [16] 2019 T1, T2, T1C, FLAIR, T1-DUAL, T2-SPIR, T1-IR, CT, PET, MRI Semantic Segmentation LeNet, ZFNet, GoogleNet, DenseNet, ResNet, AlexNet, U-Net, RNN, SVM, FCN, CRF Brain Tumor, Brain Stroke, Ischemic stroke lesion, Abdominal Organs, Intervertebral Disc
@@
6 Kar et al.  [17] 2020 Mammograms, MRI, Endoscopic OCT, Microscopic images, Radiographs, Colonoscopy, CT, X-Rays, Fundus Camera, Multi-modal, 3D Images Semantic Segmentation U-Net++, Dual U-Net, SegNet, PSPNet, U-Net, ReNet, ResNet, VGG-16, AlexNet, GoogleNet, GAN, DBN, RBM, AE, SAE, RNN, FCN Breast, Brain, Eye, Pancreas, Gland, Chest, Head, Tissues, Lungs, Cells
@@
7 Taghanaki, Saeid Asgari, et al.  [2] 2020 Microscopy, MRI, CT, EM, Fundus, X-Ray, PET Semantic Segmentation SegNet, FCN, Auto-DeepLab, RefineNet, ResNet-38, DeepLabV3 Prostate, Lung Nodule, Pancreas, Abdomen, Spleen, Breast, Heart, Spinal Cord, Retinal Vessel, Brain, Colon, Vertebral Body
@@
8 Tajbakhsh, Nima, et al.  [18] 2020 MRI, CT, X-Rays, LGE, bSSFP, DRR, Ultrasound, Angiography, FLAIR, Fundus images, Histography Semantic Segmentation, Classification, Detection Self-Learning Models, Semi-Supervised Models, Unsupervised Models, Ensembled Models Lungs, Abdomen, Breast, Tumor, Brain, Pancreas, Kidney, Liver, Heart
@@
9 Iqbal, Ahmed et al.  [19] 2021 CT, MRI, X-Ray, Mammography, Ultrasound Segmentation Vanilla-GAN, Cycle-GAN, Patch-GAN, Style-GAN, DC-GAN Retina, Brain, Skin, Heart, Breast, Bone, Lungs
@@
10 Punn and Agarwal  [20] 2021 Ultrasound, PET, X-Ray, CT, FLAIR Segmentation U-Nets’ Variants (Better U-Nets, Inception U-Nets, Attention U-Nets, Ensemble U-Nets) Heart, Lungs, Chest, Bones, Blood Vessels, Soft Tissues, Skin
@@
11 Imran et al.  [21] 2021-2022 CT, MRI, US, Microscopy, Histology, EM, PET, Colonoscopy Semantic Segmentation FCN, U-Net, Gamma-Net, SSIM, AFPNet, ResNet Brain, Chest, Breast, Skin, Prostate
@@
12 Fernando and Tsokos  [22] 2022 3D-MRI Semantic Segmentation Statistical Methods, Deep Learning Models, Hybrid Models Brain
@@
13 Xiao, Hanguang et al.  [23] 2023 MRI, CT Semantic Segmentation Different types of transformer models Abdomen, Heart, Brain, Lungs
@@
14 Wang, R. et al.  [24] 2022 X-Rays, CT, MRI, Ultrasound Semantic Segmentation U-Net, 3D-Net, RNN, ResPath, FED-Net, FCN, CDNN, GAN, Transformer models Liver, Pancreas, Heart, Lungs, Brain, Kidney, Spleen, Skin, Thyroid, Prostate, Colon
@@
15 Iqbal et al.  [25] 2023 CT, MRI, X-Rays, Echocardiograms Semantic Segmentation CNN, U-Net, DeepLabV3, ResNet-18, RF, NB, DT, LR, KNN, XGBoost Chest, Arteries, Heart
@@
16 Ilesanmi  [26] 2023 Fundus Classification, Segmentation CNN, FCNN, VGG, U-Net, LeNet-5, M-Net Arteries, Veins, Retina
@@
17 Our Survey (SLR) 2024 30 different modalities Object Detection Based segmentation Two-stage detectors, One-stage detectors, Advanced detectors Maximum Body Organs Covered
@@ 1.2 Related Surveys
Table 1 presents related surveys in the field and their comparison with our current SLR. Many surveys have been conducted related to different types of segmentation methods of biomedical images. All of them cover different types of biomedical image segmentation approaches with keeping in consideration various parameters in their scopes. As shown in Table 1, the major attractions of researchers in the last decade have been semantic segmentation using different types of machine learning models, semi-supervised learning-based models, and supervised learning-based models. A very less and insufficient attention has been paid to instance segmentation with further low attention towards object detection-based segmentation models.  Table 1 presents a comparison of all the existing surveys and distinction of our current survey. The main focus of our survey paper will be different types of object detection-based segmentation models applied for segmentation of biomedical imaging which will be the first survey conducted on this worth-attraction and important direction. Furthermore, it is evident that no proper attention has been given to the segmentation of biomedical imaging based on object detection models which are very powerful and recent techniques for the diagnosis of medical images.
Moreover, the existing surveys does not follow SLR or other standard methodology. SLR provides a structured predefined set of rules and comprehensive synthesis of research findings. It offers critical insights on a given topi, and thereby yields a well structured survey. It is unbiased as every step from searching to the final results presentation are clearly defined in an organized way.
@@ 2.4 Article’s Quality Assessment
In addition, to the above selection criteria, we excluded 17 articles based on their quality. To evaluate the quality and relevance of each article for our study, we used a quantitative metric indicating articles quality. The metric was calucalted for each article as:
@@ 3 BIOMEDICAL IMAGING MODALITIES
There are many types of biomedical images’ modalities applied for detection, diagnosis, and treatment of various types of diseases. Depending on the capturing mechanisms, materials and equipment used, and the procedure followed for their diagnosis, these images have been categorized in different classes which are all covered in this section. Various types of biomedical imaging modalities are shown in Figure 2.
@@
Although, medical radiology is a major branch of medical sciences that deals with detection, diagnosis, and treatment of various types of diseases using imaging technology, based on some standardized characterization of few imaging modalities, they have been classified under this category. The important medical radiology imaging modalities include X-rays, Computer Tomography (CT), Magnetic Resonance Imaging (MRI), and angiography images which are special types of X-rays or CT images  [28].
Ultrasound Images (US)
@@
The second important category of medical images modalities is ultrasound images which are used for diagnosis of various organs e.g., abdomen, kidneys, breast, spleen, liver, pancreas, and lungs. The important modalities of ultrasound are sonography images, color doppler, automatic breast ultrasound (ABUS), contrast enhanced ultrasound (CEUS), tissue elasticity, shear wave elastography (SWE), 3D ultrasound, and power doppler  [29].
@@
paragraphMEndoscopy and Colonoscopy images The colonoscopy and endoscopy images are used for the diagnosis and treatment of gastrointestinal tract which starts from mouth, goes through pharynx and oesophagus, stomach and small intestine, large intestine and colon. The important categories of these images include white light endoscopy (WLE), narrowband imaging (NBI), optical coherence tomography (OCT), confocal laser endomicroscopy (CLE), endoscopic ultrasound (EUS), photoacoustic endoscopy (PAE), and photoacoustic endoscope ultrasound miniprobe (PAE-UE)  [30].
Ophthalmology Imaging Modalities
@@
Ophthalmology is a branch of medical sciences that deals with the analysis, diagnosis, treatment and recovery of disorders associated with eyes. For the diagnosis and treatments of different eye parts, various types of imaging modalities are used which have different properties depending on what parts of eyes are considered and which types of diseases are targeted. The important imaging modalities used for the diagnosis and treatment of eyes include fundus photography, scanning laser ophthalmoscope (SLO), optical coherence tomography (OCT), and PAOM  [31].
Dermatology Imaging
@@
Dermatology is a branch of medical sciences that deals with the skin. The field has responsibility of both the medicinal purpose as well as surgical purpose. There are many imaging modalities that used for diagnosis and treatment of skin related diseases e.g., dermoscopy (related to skin), trichoscopy (related to scalp and hair), and onychoscopy (related to nails)  [32].
Functional Imaging
@@
MRI, CT, US, positron emission tomography (PET), and single photon emission computed tomography (SPECT) are functional images because these are used for analysing brain functionalities  [33].
Microscopic Imaging
@@
The basic concept of microscopic imaging is that it uses microscope for capturing the images of different body parts at very deep and detailed cellular levels for the detection and diagnosis of various abnormalities in the organs. There are many types of microscopic imaging modalities but important of them include selective plane illumination microscopy (SPIM), structured structure illumination microscopy (SSIM), two photon intravital microscopy (2P-IVM), two photon micro endoscopy (2P-ME)  [34].
Histopathology Imaging
@@
Histopathology is the study of the signs of the disease using the microscopic examination of a biopsy or surgical specimen that is processed and fixed onto glass slides. The images used in this process are called histopathology images. To visualize different components of the tissue under a microscope, the sections are dyed with one or more stains. The histopathology images are used for the diagnosis of different human orgnas e.g., breast  [35].
Molecular and Cellular Imaging
@@
Molecular and cellular images are very small types of images which are used for studying the finer details of structures of molecules and cells. These images are actually the microscopic images but used for the study of cells and molecules e.g., red blood celles (RBC), whie blood celles (WBC), yeast, crystal, and epithelium  [36].

### v2

@@ Abstract
Biomedical image segmentation plays a vital role in diagnosis of diseases across various organs. Deep learning-based object detection methods are commonly used for such segmentation. There exists an extensive research in this topic. However, there is no standard review on this topic. Existing surveys often lack a standardized approach or focus on broader segmentation techniques. In this paper, we conducted a systematic literature review (SLR), collected and analysed 148 articles that explore deep learning object detection methods for biomedical image segmentation. We critically analyzed these methods, identified the key challenges, and discussed the future directions. From the selected articles we extracted the results including the deep learning models, targeted imaging modalities, targeted diseases, and the metrics for the analysis of the methods. The results have been presented in tabular forms. The results are presented in three major categories including two stage detection models, one stage detection models and point-based detection models. Each article is individually analyzed along with its pros and cons. Finally, we discuss open challenges, benefits, and future directions. This SLR aims to provide the research community with a quick yet deeper understanding of these segmentation models, ultimately facilitating the development of more powerful solutions for biomedical image analysis.
Figure 1: Overview of the Research Methodology and Search Results: A step-by-step flow from article search to final selection and quality assessment, leading to the results.
@@ 1 Introduction
Biomedical image segmentation plays a vital role in the analysis and diagnosis of diseases related to various organs. In this regard, the accurate identification of the region of interest in the biomedical images is a significant phase for identifying affected regions in the image [1]. There are several alternative methods to achieve this segmentation. object detection based methods is a type of instance segmentation that is very common in this domain. Deep learning-based object detection methods [2, 3] offer significant advantages over the other alternatives. Considering the prominence of deep learning based object detection-based methods for biomedical image segmentation, we believe a comprehensive review of this field can be a a valuable work to provide a quick insight into the field. Although, there are few reviews (section 1.2) in the related topic, there is no standard review on the target direction.
@@ 1.1 Background and Motivation
In the instance segmentation, each pixel of image is assigned to some specific instance which has a significant role in biomedical imaging for the identification and analysis of various types of diseases or same type of disease with different locations in the image. Different types of algorithms are used for instance segmentation e.g., supervised learning (SL) models and semi-supervised learning (SSL) approaches. Each approach has its own drawbacks and benefits associated with them. Recently, fully supervised models have got proper attention for biomedical image instance segmentation which are based on human experts’ knowledge for annotations. These models are very laborious and expensive in terms of obtaining dense and accurate annotations  [12]. To overcome the problem of time consumption and human efforts associated with annotation for supervised learning approaches, some semi-supervised models have been proposed which require small set of annotated biomedical image samples as compared to the fully supervised instance segmentation models  [13]. The instance segmentation is mainly categorized into two main types namely proposal-free  [14] and proposal-based instance segmentation  [15]. The former category is based on morphological properties and instance-aware features whereas the latter is type is based on object detection approach which is further divided into two stage detectors, one stage detectors, and point-based detectors or advanced detectors.
@@
Table 1: Related Surveys and Comparison with Our SLR.
@@
S.No Review ID Covered up to Year Imaging Modalities Processing Tasks Performed Models Targeted Diseases/Organs Covered
@@
1 R01 [11] 2016 MRI (T1, T1-Gd, T2, FLAIR) Automatic, Semantic Segmentation CNN, ML, Generative Models Brain Tumor
@@
2 R02 [16] 2018 CT, MRI, PET, X-Rays, Ultrasound Segmentation, Classification, Diagnosis CNN different variants + ML models Lungs, Breast, Brain, Thyroid
@@
3 R03 [17] 2019 Clinical, X-Rays, CT, MRI, Ultrasound, OCT, Microscopic Images Semantic Segmentation, Classification, Analysis ML, CNN, RBM, Auto encoders, GANs, RNN, FCN, U-Net, ResNet, V-Net Brain, Breast, Liver, Spleen, Heart, Stomach, Ventricles, Arteries, Skin
@@
4 R04 [18] 2019 PET, CT, X-Ray, Ultrasound, Histological Images Classification, Segmentation VGG-16, GoogleNet, AlexNet, ResNet, ResUnet, SegNet, UNet++, RD-Unet, RNN, ML Models Not Mentioned
@@
5 R05 [19] 2019 T1, T2, T1C, FLAIR, T1-DUAL, T2-SPIR, T1-IR, CT, PET, MRI Semantic Segmentation LeNet, ZFNet, GoogleNet, DenseNet, ResNet, AlexNet, U-Net, RNN, SVM, FCN, CRF Brain Tumor, Brain Stroke, Ischemic stroke lesion, Abdominal Organs, Intervertebral Disc
@@
6 R06 [20] 2020 Mammograms, MRI, Endoscopic OCT, Microscopic images, Radiographs, Colonoscopy, CT, X-Rays, Fundus Camera, Multi-modal, 3D Images Semantic Segmentation U-Net++, Dual U-Net, SegNet, PSPNet, U-Net, ReNet, ResNet, VGG-16, AlexNet, GoogleNet, GAN, DBN, RBM, AE, SAE, RNN, FCN Breast, Brain, Eye, Pancreas, Gland, Chest, Head, Tissues, Lungs, Cells
@@
7 R07 [5] 2020 Microscopy, MRI, CT, EM, Fundus, X-Ray, PET Semantic Segmentation SegNet, FCN, Auto-DeepLab, RefineNet, ResNet-38, DeepLabV3 Prostate, Lung Nodule, Pancreas, Abdomen, Spleen, Breast, Heart, Spinal Cord, Retinal Vessel, Brain, Colon, Vertebral Body
@@
8 R08 [21] 2020 MRI, CT, X-Rays, LGE, bSSFP, DRR, Ultrasound, Angiography, FLAIR, Fundus images, Histography Semantic Segmentation, Classification, Detection Self-Learning Models, Semi-Supervised Models, Unsupervised Models, Ensembled Models Lungs, Abdomen, Breast, Tumor, Brain, Pancreas, Kidney, Liver, Heart
@@
9 R09 [22] 2021 CT, MRI, X-Ray, Mammography, Ultrasound Segmentation Vanilla-GAN, Cycle-GAN, Patch-GAN, Style-GAN, DC-GAN Retina, Brain, Skin, Heart, Breast, Bone, Lungs
@@
10 R10 [23] 2021 Ultrasound, PET, X-Ray, CT, FLAIR Segmentation U-Nets’ Variants (Better U-Nets, Inception U-Nets, Attention U-Nets, Ensemble U-Nets) Heart, Lungs, Chest, Bones, Blood Vessels, Soft Tissues, Skin
@@
11 R11 [24] 2021-2022 CT, MRI, US, Microscopy, Histology, EM, PET, Colonoscopy Semantic Segmentation FCN, U-Net, Gamma-Net, SSIM, AFPNet, ResNet Brain, Chest, Breast, Skin, Prostate
@@
12 R12 [25] 2022 3D-MRI Semantic Segmentation Statistical Methods, Deep Learning Models, Hybrid Models Brain
@@
13 R13 [26] 2023 MRI, CT Semantic Segmentation Different types of transformer models Abdomen, Heart, Brain, Lungs
@@
14 R14 [27] 2022 X-Rays, CT, MRI, Ultrasound Semantic Segmentation U-Net, 3D-Net, RNN, ResPath, FED-Net, FCN, CDNN, GAN, Transformer models Liver, Pancreas, Heart, Lungs, Brain, Kidney, Spleen, Skin, Thyroid, Prostate, Colon
@@
15 R15 [28] 2023 CT, MRI, X-Rays, Echocardiograms Semantic Segmentation CNN, U-Net, DeepLabV3, ResNet-18, RF, NB, DT, LR, KNN, XGBoost Chest, Arteries, Heart
@@
16 R16 [29] 2023 Fundus Classification, Segmentation CNN, FCNN, VGG, U-Net, LeNet-5, M-Net Arteries, Veins, Retina
@@
17 Ours 2024 30 different modalities Object Detection Based segmentation Two-stage detectors, One-stage detectors, Advanced detectors Maximum Body Organs Covered
@@ 1.2 Related Surveys
Table 1 presents related surveys in the field and their comparison with our current SLR. Many surveys have been conducted related to different types of segmentation methods of biomedical images. All of them cover different types of biomedical image segmentation approaches with keeping in consideration various parameters in their scopes. As shown in Table 1, the major attractions of researchers in the last decade have been semantic segmentation using different types of machine learning models, semi-supervised learning-based models, and supervised learning-based models. A very less and insufficient attention has been paid to instance segmentation with further low attention towards object detection-based segmentation models.  Table 1 presents a comparison of all the existing surveys and distinction of our current survey. The main focus of our survey paper will be different types of object detection-based segmentation models applied for segmentation of biomedical imaging which will be the first survey conducted on this worth-attraction and important direction. Furthermore, it is evident that no proper attention has been given to the segmentation of biomedical imaging based on object detection models which are very powerful and recent techniques for the diagnosis of medical images. One similar survey has been conducted by Gui [30]; however, their focus is on remote sensing, and they did not follow the Systematic Literature Review (SLR) methodology.
The lack of following the SLR or other standard methodologies has been consistent across all existing review articles in the field. This review will be among the pioneer articles in AI-based image processing, particularly in image segmentation. SLR provides a structured predefined set of rules and comprehensive synthesis of research findings. It offers critical insights on a given topi, and thereby yields a well structured survey. It is unbiased as every step from searching to the final results presentation are clearly defined in an organized way.
@@ 2.4 Article’s Quality Assessment
In addition, to the above selection criteria, we excluded 15 articles based on their quality. To evaluate the quality and relevance of each article for our study, we used a quantitative metric indicating articles quality. The metric was calucalted for each article as:
@@ 3 BIOMEDICAL IMAGING MODALITIES
There are many types of biomedical images’ modalities applied for detection, diagnosis, and treatment of various types of diseases. Depending on the capturing mechanisms, materials and equipment used, and the procedure followed for their diagnosis, these images have been categorized in different classes which are all covered in this section. Various types of biomedical imaging modalities are shown in Figure 2, and described below.
@@
1) Medical Radiology Images: Although, medical radiology is a major branch of medical sciences that deals with detection, diagnosis, and treatment of various types of diseases using imaging technology, based on some standardized characterization of few imaging modalities, they have been classified under this category. The important medical radiology imaging modalities include X-rays, Computer Tomography (CT), Magnetic Resonance Imaging (MRI), and angiography images which are special types of X-rays or CT images  [32].
@@
2) Ultrasound Images (US): The second important category of medical images modalities is ultrasound images which are used for diagnosis of various organs e.g., abdomen, kidneys, breast, spleen, liver, pancreas, and lungs. The important modalities of ultrasound are sonography images, color doppler, automatic breast ultrasound (ABUS), contrast enhanced ultrasound (CEUS), tissue elasticity, shear wave elastography, 3D ultrasound, and power doppler  [33].
@@
3) Endoscopy and Colonoscopy Images: The colonoscopy and endoscopy images are used for the diagnosis and treatment of gastrointestinal tract which starts from mouth, goes through pharynx and oesophagus, stomach and small intestine, large intestine and colon. The important categories of these images include white light endoscopy (WLE), narrowband imaging (NBI), optical coherence tomography (OCT), confocal laser endomicroscopy (CLE), endoscopic ultrasound (EUS), photoacoustic endoscopy (PAE), and photoacoustic endoscope ultrasound miniprobe (PAE-UE)  [34].
@@
4) Ophthalmology Imaging Modalities: Ophthalmology is a branch of medical sciences that deals with the analysis, diagnosis, treatment and recovery of disorders associated with eyes. For the diagnosis and treatments of different eye parts, various types of imaging modalities are used which have different properties depending on what parts of eyes are considered and which types of diseases are targeted. The important imaging modalities used for the diagnosis and treatment of eyes include fundus photography, scanning laser ophthalmoscope (SLO), optical coherence tomography (OCT), and PAOM  [35].
@@
5) Dermatology Imaging: Dermatology is a branch of medical sciences that deals with the skin. The field has responsibility of both the medicinal purpose as well as surgical purpose. There are many imaging modalities that used for diagnosis and treatment of skin related diseases e.g., dermoscopy (related to skin), trichoscopy (related to scalp and hair), and onychoscopy (related to nails)  [36].
@@
6) Functional Imaging: MRI, CT, US, positron emission tomography (PET), and single photon emission computed tomography (SPECT) are functional images because these are used for analysing brain functionalities  [37].
@@
7) Microscopic Imaging: The basic concept of microscopic imaging is that it uses microscope for capturing the images of different body parts at very deep and detailed cellular levels for the detection and diagnosis of various abnormalities in the organs. There are many types of microscopic imaging modalities but important of them include selective plane illumination microscopy (SPIM), structured structure illumination microscopy (SSIM), two photon intravital microscopy (2P-IVM), two photon micro endoscopy (2P-ME)  [38].
@@
8) Histopathology Imaging: Histopathology is the study of the signs of the disease using the microscopic examination of a biopsy or surgical specimen that is processed and fixed onto glass slides. The images used in this process are called histopathology images. To visualize different components of the tissue under a microscope, the sections are dyed with one or more stains. The histopathology images are used for the diagnosis of different human orgnas e.g., breast  [39].
@@
9) Molecular and Cellular Imaging Molecular and cellular images are very small types of images which are used for studying the finer details of structures of molecules and cells. These images are actually the microscopic images but used for the study of cells and molecules e.g., red blood celles (RBC), whie blood celles (WBC), yeast, crystal, and epithelium  [40].